AN ANALYSIS OF VIBRATION-TORSION-ROTATION INTERACTIONS AND SPECTRA FOR A MOLECULE WITH AN INTERNAL ROTOR

\underline{YUN-BO DUAN}, Department of Electronics and Computer Science, Toyama University, Toyama 930, Japan; KOJIRO TAKAGI, Department of Physics, Toyama University, Toyama 930, Japan.

Based on the analysis of order of magnitude of the term H_{mln} in a Hamiltonian, a new scheme of the ordering by magnitude is proposed to analyze the vibration-torsion-rotation interactions for molecules with internal rotations, in which the order of the torsional angular momentum term with P_{γ}^{2} is considered to be just between those of the vibrational angular momentum term with p_{k}^{2} and the rotational angular momentum term with P^{2} (or P_{a}^{2}). Therefore, the importance of torsional problem is emphasized. The sequential contact transformation technique is applied to the vibration-torsion-rotation Hamiltonian for a molecule containing a threefold symmetric internal rotor, which makes it possible to analyze clearly various higherorder vibration-torsion-rotation interactions and their effects in the vibration-torsion-rotation spectra. A reduced torsionrotation Hamiltonian for the molecule with C_{3v}(M) symmetry is given in the new ordering scheme and is applied to analyze the spectra of CH_{3}OH and CH_{3}^{18}OH. Some new lines are observed for CH_{3}^{18}OH.